OpenStudy (anonymous):

The formula to find the perimeter of a rectangle is: P = 2W + 2L. To solve for " W " in this formula, which steps should be followed? A.) Subtract 2L from both sides of the equation, then subtract 2 from both sides of the equation. B.) Subtract 2L from both sides of the equation, then divide both sides of the equation by 2. C.) Divide both sides of the equation by 2L, then subtract 2 from both sides of the equation. D.) Divide both sides of the equation by 2L +

OpenStudy (anonymous):

my answer is C

OpenStudy (anonymous):

It's B

OpenStudy (anonymous):

P = 2W + 2L -2L P-2L=2W Divide by 2 1/2P-L=W

OpenStudy (hexagon001):

P=2W +2L (SUBTRACT 2L FROM BOTH SIDES) P-2L=2W+2L-2L P-2L=2W (NOW DIVIDE BY 2 TO GET THE 'W' BY ITSELF) (P-2L)/2 =W SO W= (P-2L)/2

OpenStudy (hexagon001):

sorry for the caps lock

OpenStudy (anonymous):

Thank you for the explanation both @RoseDryer and @hexagon001